# Zero-Shot Referring Expression Comprehension via Visual-Language True/False Verification

### v1

@@ Abstract
Index Terms:
Referring expression comprehension, large language modeling inference, zero-shot reasoning
@@ I Introduction
Despite strong progress in large vision–language models (VLMs), e.g., CLIP[2], LLaVA[3], GPT-4o[4], REC remains difficult. The task demands fine-grained language understanding and grounded visual reasoning to select one instance among many look-alike objects. Descriptions blend attributes, parts, and spatial/ordinal cues, often in scenes with occlusion and scale changes, so that close similarities are common. Furthermore, off-the-shelf VLMs without REC training rarely produce accurate bounding boxes as they are not calibrated for instance-level localization.
@@
On RefCOCO[8], RefCOCO+, and RefCOCOg[9], our verification-first workflow achieves ACC@0.5 scores of 81.2%, 76.3%, and 75.9%, respectively. This marks a substantial improvement over the zero-shot GroundingDINO baseline without RefCOCO training (50.4%, 51.4%, 60.4%). Our approach sets a new state of the art for these benchmarks under zero-shot conditions. Here, ‘zero-shot’ means that no model is trained on the REC task, the detector is COCO-clean (trained without COCO[7] data), and the VLM is used strictly off-the-shelf—though its pretraining corpus may include COCO images, which is not explicitly known. It also surpasses the reported performance of GroundingDINO with RefCOCO training by 11.3% on average. Compared with a single-shot selection approach that uses hand-crafted prompts over the same proposals and VLMs, our proposed True/False verification workflow improves ACC@0.5(%) by 18.0% on average across the three benchmarks and two VLMs.
@@ II-B Verification-First Workflow
Zero-shot VLM (baseline) GPT-4o (organic / vanilla selection) 8.4 7.3 9.1
@@
Zero-shot on REC (ours) Selection (GPT-4o, single-shot) 70.1 73.5 65.4 66.7 68.3 60.2 69.7 68.4
Zero-shot on REC (ours) Selection (GPT-4o, majority voting) 71.7 67.4 69.9
Zero-shot on REC (ours) Verification-first (GPT-4o) 81.2 86.2 74.5 76.3 81.2 69.4 75.9 74.6
@@ IV Experiments
  • GPT-4o (vanilla REC). Off-the-shelf GPT-4o is shown the image with description and asked to generate the best-matching bounding box in a single prompt (no REC training).
@@
  • Verification-first (GPT-4o, ours). Same verification pipeline as above but using GPT-4o instead of LLaVA.
@@
  • Selection (GPT-4o, single-shot, ours). Same selection pipeline as above but using GPT-4o instead of LLaVA.
  • Selection (GPT-4o, majority voting, ours). Same selection pipeline as above, using GPT-4o, but using majority voting of three runs. Temperature is configured to 1.0.
We use YOLO-World as a basic class-conditioned detector (no grounding), trained without COCO for dataset cleanliness. This COCO-clean choice comes with a modest cost, i.e., approximately 10 percentage points mAP@0.5 lower than comparable COCO-trained YOLO variants on detection tasks. For VLMs, we evaluate GPT-4o and LLaVA-vicuna-13b off-the-shelf. A “organic” GPT-4o baseline attains less than 10% ACC@0.5 on these benchmarks, suggesting that it is not REC-trained. The strong numbers we obtain stem from the workflow rather than hidden task-specific pretraining.
As shown in Table 2, except for the fully supervised CogVLM, our verification-first (GPT-4o) approach outperforms all other approaches across RefCOCO/RefCOCO+/RefCOCOg. Using the same detector, VLM, and proposal sets, verification consistently beats selection, including the selection with majority voting, by about 6 to 10 absolute points across most of the splits, underscoring that workflow—not just model choice—drives the gains. We also observe that GPT-4o surpasses LLaVA in both verification and selection, indicating that VLM capability is critical in zero-shot REC. In summary: (i) a carefully designed workflow, like the True/False verification workflow, can deliver large improvements even with identical components, and (ii) fundamental models (a basic, non-grounding detector and a general-purpose VLM) can be composed to solve a composite task like REC without task-specific training—though the strength of each component still matters.

### v3

# Zero-Shot Referring Expression Comprehension via Vision-Language True/False Verification

@@ I Introduction
Despite strong progress in large vision–language models (VLMs), e.g., CLIP[2], LLaVA[3], GPT-5[15], REC remains difficult. The task demands fine-grained language understanding and grounded visual reasoning to select one instance among many look-alike objects. Descriptions blend attributes, parts, and spatial/ordinal cues, often in scenes with occlusion and scale changes, so that close similarities are common. Furthermore, off-the-shelf VLMs without REC training rarely produce accurate bounding boxes as they are not calibrated for instance-level localization.
@@
On RefCOCO[8], RefCOCO+, and RefCOCOg[9], our verification-first workflow with GPT-5 achieves ACC@0.5 scores of 79.3%, 74.2%, and 72.4%, respectively. This marks a substantial improvement over the zero-shot GroundingDINO baseline without RefCOCO training (50.4%, 51.4%, 60.4%). Our approach sets a new state of the art for these benchmarks under zero-shot conditions. Here, ‘zero-shot’ means that no model is trained on the REC task, the detector is COCO-clean (trained without COCO[7] data), and the VLM is used strictly off-the-shelf—though its pretraining corpus may include COCO images, which is not explicitly known. It also surpasses the reported performance of GroundingDINO with RefCOCO training by 7.2% on average. Compared with a single-shot selection approach that uses hand-crafted prompts over the same proposals and VLMs, our proposed True/False verification workflow improves ACC@0.5(%) by 12.3% on average across the three benchmarks and two VLMs.
@@ II-B Verification-First Workflow
Zero-shot VLM (baseline) GPT-5 (organic / vanilla selection) 11.7 10.1 12.5
@@
Zero-shot on REC (ours) Selection (GPT-5, single-shot) 70.1 73.5 65.4 66.7 68.3 60.2 69.7 68.4
Zero-shot on REC (ours) Selection (GPT-5, majority voting) 71.7 67.4 69.9
Zero-shot on REC (ours) Binary verification (GPT-5) 79.3 85.6 70.4 74.2 80.4 65.2 72.4 71.5
@@ IV Experiments
  • GPT-5 (vanilla REC). Off-the-shelf GPT-5 is shown the image with description and asked to generate the best-matching bounding box in a single prompt (no REC training).
@@
  • Verification-first (GPT-5, ours). Same verification pipeline as above but using GPT-5 instead of LLaVA.
@@
  • Selection (GPT-5, single-shot, ours). Same selection pipeline as above but using GPT-5 instead of LLaVA.
  • Selection (GPT-5, majority voting, ours). Same selection pipeline as above, using GPT-5, but using majority voting of three runs. Temperature is configured to 1.0.
We use YOLO-World as a basic class-conditioned detector (no grounding), trained without COCO for dataset cleanliness. This COCO-clean choice comes with a modest cost, i.e., approximately 10 percentage points mAP@0.5 lower than comparable COCO-trained YOLO variants on detection tasks. For VLMs, we evaluate GPT-5 and LLaVA-vicuna-13b off-the-shelf. A “organic” GPT-5 baseline attains less than 15% ACC@0.5 on these benchmarks, suggesting that it is not REC-trained. The strong numbers we obtain stem from the workflow rather than hidden task-specific pretraining.
As shown in Table 2, except for the fully supervised CogVLM, our verification-first (GPT-5) approach outperforms most other approaches across RefCOCO/RefCOCO+/RefCOCOg. Using the same detector, VLM, and proposal sets, verification consistently beats selection, including the selection with majority voting, by about 5 to 10 absolute points across most of the splits, underscoring that workflow—not just model choice—drives the gains. We also observe that GPT-5 surpasses LLaVA in both verification and selection, indicating that VLM capability is critical in zero-shot REC. In summary: (i) a carefully designed workflow, like the True/False verification workflow, can deliver large improvements even with identical components, and (ii) fundamental models (a basic, non-grounding detector and a general-purpose VLM) can be composed to solve a composite task like REC without task-specific training—though the strength of each component still matters.